OpenStudy (anonymous):

solve by completing the square x^2-16x= 10

OpenStudy (anonymous):

take the -16 term divide it by 2 and square it then add it to both sides of the equation: \[\frac{-16}{2}= -8\rightarrow (-8)^2=64\] so the equation turns into: \[x^2-16x+64=10+64\] you can factor the left hand side: \[(x-8)(x-8)=74 \rightarrow (x-8)^2=74\rightarrow (x-8)^2-74=0\]

OpenStudy (anonymous):

\[(x-8)^2=74\]

OpenStudy (anonymous):

so x would = 74?

OpenStudy (anonymous):

no x=8

OpenStudy (anonymous):

y=-74

OpenStudy (anonymous):

vertex: ( 8,-74)

OpenStudy (anonymous):

are you any good with graphing?

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

i have one i have posted thats been giving me trouble

OpenStudy (anonymous):

the answer for this one was in radical form which it did not state, ty for the help you still get a medal

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

nadeem solve this call vertex from

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

\[(x-8)^2=10+64\]\[x-8=\pm \sqrt{74}\]\[x=8\pm \sqrt{74}\] vertex from: \[x^2-16x-10\]\[1(x-8)^2-10-(1*64)\]\[(x-8)^2-10-64\]\[y=(x-8)^2-74\] vertex=(8,-74)